# Fundamental thresholds for computational and erasure errors via the coherent information

### v1

@@ Abstract
Quantum error correcting (QEC) codes protect quantum information against environmental noise. Computational errors caused by the environment change the quantum state within the qubit subspace, whereas quantum erasures correspond to the loss of qubits at known positions. Correcting either type of error involves different correction mechanisms, which makes studying the interplay between erasure and computational errors particularly challenging. In this work, we propose a framework based on the coherent information (CI) of the mixed-state density operator associated to noisy QEC codes, for treating both types of errors together. We show how to rigorously derive different families of statistical mechanics mappings for generic stabilizer QEC codes in the presence of both types of errors. We observe that the erasure errors enter as a classical average over fully depolarizing channels. Further, we show that computing the CI for erasure errors only can be done efficiently upon sampling over erasure configurations. We then test our approach on the 2D toric and color codes and compute optimal thresholds for erasure errors only, finding a $50\%$ threshold for both codes. This strengthens the notion that both codes share the same optimal thresholds. When considering both computational and erasure errors, the CI of small-size codes yields thresholds in very accurate agreement with established results that have been obtained in the thermodynamic limit. We thereby further establish the CI as a practical tool for studying optimal thresholds under realistic noise and as a means for uncovering new relations between QEC codes and statistical physics models.
@@ I Introduction
Storing and manipulating quantum information in noisy quantum devices is one of the main challenges in the field of quantum technologies. In the framework of quantum computing, quantum error correction (QEC) Lidar and Brun (2013); Terhal (2015) is the main technique for protecting quantum information from noise in a scalable manner and is the key element on the route towards fault-tolerant quantum computing. QEC works by encoding one or several logical qubits into many noisy physical qubits and using the redundant degrees of freedom for the detection and subsequent removal of errors. This allows one to reduce the overall error rate of the logical qubit(s) compared to the bare physical qubits, provided the error rates of the faulty operations needed to operate the QEC code fall below code- and noise-model dependent critical threshold values. There has been impressive recent progress in realizing error-corrected and fault-tolerantly operated logical qubits in a variety of physical platforms, including superconducting circuits AI (2024); Google Quantum AI (2023, 2021); Krinner et al. (2022); Zhao et al. (2022); Andersen et al. (2020); Sivak et al. (2023); Gupta et al. (2024); Hetényi and Wootton (2024); Lacroix et al. (2024), trapped ions Postler et al. (2022); Ryan-Anderson et al. (2022); da Silva et al. (2024); Berthusen et al. (2024); Ryan-Anderson et al. (2021, 2024); Pogorelov et al. (2024); Postler et al. (2024); Huang et al. (2024a), and neutral atoms Bluvstein et al. (2024); Bedalov et al. (2024); Reichardt et al. (2024); Rodriguez et al. (2024). Fault-tolerant operation of logical qubits allows one to scale up the underlying QEC codes to larger distances, and thereby systematically suppress associated logical error rates. This suppression works as long as the failure rate of all physical components is smaller than a certain threshold error rate Aharonov and Ben-Or (2008); Knill et al. (1998); Kitaev (2003); Shor (1996). Recent experiments have demonstrated the suppression of logical error rates expected below threshold AI (2024); Google Quantum AI (2023); da Silva et al. (2024); Gupta et al. (2024); Hong et al. (2024); Sivak et al. (2023).
@@
The exact threshold value depends on the QEC code and decoding procedure used for recovering the logical information. Among all decoding strategies, Maximum Likelihood Decoding (MLD) stands out because it has the highest possible threshold, dubbed optimal threshold. In principle, obtaining optimal thresholds requires to simulate QEC by sampling over all possible error processes and then computing the corrective operation using MLD. The latter requires, in general, to solve an exponentially hard problem Iyer and Poulin (2015); Fuentes et al. (2021) every time an error is corrected, thus severely limiting the size of the codes one can simulate. However, recently it has been pointed out that the coherent information (CI) Schumacher and Nielsen (1996) of the mixed state associated to the noisy qubits forming the QEC code shows a discontinuity at the optimal threshold Fan et al. (2024); Colmenarez et al. (2024). In fact, the CI captures the information left in the noisy mixed-state that can, in principle, be recovered Schumacher and Nielsen (1996). Therefore, the CI provides a practical, generic tool for estimating optimal thresholds without resorting to numerical simulation of noisy QEC cycles and performing MLD.
@@
Once one or more physical qubits are lost or erased, the corrective operation is equivalent to trying to retrieve the logical qubit from the remaining physical qubits. Several QEC schemes for the correction of erasure errors have been designed Grassl et al. (1997); Amaro et al. (2020); Auger et al. (2017); Delfosse and Nickerson (2021); Goto and Uchikawa (2014); McLauchlan et al. (2024); Morley-Short et al. (2019); Nagayama et al. (2017); Siegel et al. (2023); Steinberg et al. (2024); Vodola et al. (2022), including first experimental demonstrations of deterministic correction of qubit loss Stricker et al. (2020); Reichardt et al. (2024). Let us note that optimal decoding of QEC codes in the presence of qubit loss is a fundamentally different problem compared to computational errors. For instance, in the case of the surface code, there is a linear-time maximum likelihood decoder (MLD) under qubit loss Delfosse and Zémor (2020), whereas efficient optimal decoding for computational errors is only possible for uncorrelated bit and phase-flip noise Bravyi et al. (2014). Another difference concerns the minimum number of physical qubits required to correct a single error on a qubit, which is four for qubit losses Grassl et al. (1997) and five for an arbitrary computational error Laflamme et al. (1996). Furthermore, optimal thresholds of QEC codes under erasure errors are usually higher than for computational errors Stace et al. (2009). In fact, given the positive trade-off in correcting qubit losses instead of computational errors, it has been proposed that erasure conversion, i.e., converting computational errors into qubit losses, could significantly improve the overall performance of error correction protocols Wu et al. (2022); Kang et al. (2023); Sahay et al. (2023); Quinn et al. (2024).
@@
On the other hand, we numerically compute the CI for different values of the error probability and find that the finite-size crossings of the CI in small distance codes yield very accurate approximations to the optimal error thresholds compared to rigorous solutions of the corresponding statistical mechanics models Stace et al. (2009); Stace and Barrett (2010); Ohzeki (2012). This provides further evidence of the utility of the CI in estimating optimal thresholds of QEC codes from small code instances Colmenarez et al. (2024). Let us note that in this work we do not intend to solve the statistical mechanics models and pinpoint their phase transition, as done in several other works Dennis et al. (2002); Chubb and Flammia (2021). Instead, we use the models as a mean to numerically compute the CI and extract optimal thresholds directly from it.
The structure of the paper is as follows. In Sec. II a summary of the main results of this work is presented. In Sec. III the main concepts used are explained in detail. In Sec. IV we present the main derivations of our work in detail. In Sec. V we show the numerical calculations of CI for surface and color codes. Finally, concluding remarks are presented in Sec. VI.
@@ II.2 Coherent information mapping of CSS codes under depolarizing noise
Table 1: Pauli error probabilities for depolarizing channel of Eq. (3) on each site $\ell$ in term of the random variables $\eta^{x}_{\ell}$ and $\eta^{z}_{\ell}$. $C^{(1)}$ shows the connection to the error chains in Sec. IV.
@@
$H[\{\eta^{x}\},\{\eta^{Z}\}]=\sum_{\ell}\eta^{x}_{\ell}\left(J_{x}-\dfrac{J_{1}}{2}\right)P^{X}_{\ell}+$
$\eta^{z}_{\ell}\left(J_{z}-\dfrac{J_{1}}{2}\right)P^{Z}_{\ell}+\eta^{z}_{\ell}\eta^{x}_{\ell}J_{1}P^{Z}_{\ell}P^{X}_{\ell}.$ (5)
@@ II.3 Coherent information mapping of CSS codes under depolarizing and erasure errors
The last equation shows the processes that reduce the CI: (i) qubit losses can degrade the logical qubit to a classical bit or completely destroy it, and (ii) the cost of computational errors and their interplay with qubit losses is modeled by the partition functions $\mathcal{Z}_{D_{l}}$, with missing bonds at the erased qubit positions. We can reason about these two processes as follows: erasure errors weaken the ability to protect quantum information by rendering some stabilizers ill-defined and effectively reducing the number of degrees of freedom used for erorr detection and correction. However, in the absence of computational errors, this does not affect the logical qubit unless at least one logical operator becomes ill-defined. In the presence of computational errors, there are fewer operators that commute with the well-defined stabilizers, but the partition functions $\mathcal{Z}_{D,l}$ still count the number of operators that commute with the well-defined stabilizers. Note that $D_{l}$ now denotes the set of operators that anti-commute with the well-defined logical operators. For details, see Sec. IV.3.
@@ III.1 Stabilizer codes
In particular we are interested in codes for which the stabilizer generators $g_{i}$ are composed of either $X$ or $Z$ Pauli operators, i.e. Calderbank-Shor-Steane (CSS) codes Calderbank and Shor (1996); Steane (1997). This class of codes allows the logical operators $\{O^{x}_{i},O^{z}_{i}\}$ to be made of only $Z$ or $X$ Pauli operators. Within the class of CSS codes we are interested in topological codes Kitaev (2006). They are defined on $D-$dimensional lattices and encode a fixed number $k$ of logical qubits, i.e. $k$ is independent of $n$ and only dependent on the topological properties of the manifold the lattice is embedded in. The stabilizer generators $\{g_{i}\}$ are defined on unit cells of the lattice and each qubit is shared by a constant number of stabilizer generators. In this work we focus on two-dimensional toric/surface and color code Kitaev (1997); Dennis et al. (2002); Bombin and Martin-Delgado (2007a) which are amongst today’s leading contenders for the realization of scalable fault-tolerant quantum computing AI (2024); Google Quantum AI (2023, 2021); Krinner et al. (2022); Zhao et al. (2022); Andersen et al. (2020); Postler et al. (2022); Ryan-Anderson et al. (2022); da Silva et al. (2024); Berthusen et al. (2024); Ryan-Anderson et al. (2021, 2024); Pogorelov et al. (2024); Postler et al. (2024); Huang et al. (2024a); Bluvstein et al. (2024); Lacroix et al. (2024); Rodriguez et al. (2024). We refer to Appendix B for definitions and details of the topological codes studied in this work.
@@ III.2.1 Erasure channel
There is a probability $e$ that the state $\rho$ is replaced by a maximally mixed state. We are interested in the case when we know the qubit is lost due to the action of the channel. Therefore, a more convenient way to describe this process is as a classical ensemble of states $\mathcal{E}_{n}(\rho)=\{(P_{i},\rho_{i})\}$, where $P_{i}$ is the probability of recording the state $n-$qubit state $\rho_{i}$ after applying the channel $\mathcal{E}_{n}$. In the case of a single-qubit erasure channel, we have
$\mathcal{E}_{1}(\rho)=\{(1-e,\rho),(e,\mathbb{1}/2)\}.$ (13)
For simplicity we work directly with the classical ensemble of states and carry out the derivation of the mappings replacing the erased state by a maximally mixed state. As will be discussed in Sec. IV.1, the erased degrees of freedom do not contribute to the CI. Now we generalize Eq. (13) to $n-$qubit states:
$\mathcal{E}_{n}(\rho)=\{(P(l),\frac{1}{2^{m}}\mathbb{1}_{A}\otimes\operatorname{Tr}_{A_{l}}(\rho))\}.$ (14)
Here $A_{l}$ denotes the set of erased qubits and $l=1,2,..,2^{n}$ runs over all possible configurations of erased qubits. $m\equiv m(l)$ is the number of qubits in the erased region $A_{l}$. When erasures are independent on each qubit, e.g. all qubits experience the same erasure operation given by Eq. (13) with the same probability $e$, then $P(l)=(1-e)^{1-m}e^{m}$. In general a distance-$d$ QEC code can correct all erasure configurations with $m<d$. All erasure configurations for which at least one representative of each logical operator remains well-defined is fully correctable.

### v2

@@ Abstract
Quantum error correcting (QEC) codes protect quantum information against environmental noise. Computational errors caused by the environment change the quantum state within the qubit subspace, whereas quantum erasures correspond to the loss of qubits at known positions. Correcting either type of error involves different correction mechanisms, which makes studying the interplay between erasure and computational errors particularly challenging. In this work, we propose a framework based on the coherent information (CI) of the mixed-state density operator associated to noisy QEC codes, for treating both types of errors together. We show how to rigorously derive different families of statistical mechanics mappings for generic stabilizer QEC codes in the presence of both types of errors. We observe that the erasure errors enter as a classical average over fully depolarizing channels. Further, we show that computing the CI for erasure errors only can be done efficiently upon sampling over erasure configurations. We then test our approach on the 2D toric and color codes and compute optimal thresholds for erasure errors only, finding a $50\%$ threshold for both codes. This strengthens the notion that both codes share the same optimal thresholds. When considering both computational and erasure errors, the CI of small-size codes yields thresholds in very accurate agreement with established results that have been obtained in the thermodynamic limit. Next, we perform a similar analysis for a low-density parity-check (LDPC) code — the lift-connected surface code. We find a $50\%$ threshold under erasure errors alone and, for the first time, derive the exact statistical mechanics mappings in the presence of both computational and erasure errors. We thereby further establish the CI as a practical tool for studying optimal thresholds for code classes beyond topological codes under realistic noise, and as a means for uncovering new relations between QEC codes and statistical physics models.
@@ I Introduction
Storing and manipulating quantum information in noisy quantum devices is one of the main challenges in the field of quantum technologies. In the framework of quantum computing, quantum error correction (QEC) Lidar and Brun (2013); Terhal (2015) is the main technique for protecting quantum information from noise in a scalable manner and is the key element on the route towards fault-tolerant quantum computing. QEC works by encoding one or several logical qubits into many noisy physical qubits and using the redundant degrees of freedom for the detection and subsequent removal of errors. This allows one to reduce the overall error rate of the logical qubit(s) compared to the bare physical qubits, provided the error rates of the faulty operations needed to operate the QEC code fall below code- and noise-model dependent critical threshold values Aharonov and Ben-Or (2008); Knill et al. (1998); Kitaev (2003); Shor (1996). . There has been impressive recent progress in realizing error-corrected and fault-tolerantly operated logical qubits in a variety of physical platforms, including superconducting circuits AI (2024); Google Quantum AI (2023, 2021); Krinner et al. (2022); Zhao et al. (2022); Andersen et al. (2020); Sivak et al. (2023); Gupta et al. (2024); Hetényi and Wootton (2024); Lacroix et al. (2024), trapped ions Postler et al. (2022); Ryan-Anderson et al. (2022); da Silva et al. (2024); Berthusen et al. (2024); Ryan-Anderson et al. (2021, 2024); Pogorelov et al. (2024); Postler et al. (2024); Huang et al. (2024a), and neutral atoms Bluvstein et al. (2024); Bedalov et al. (2024); Reichardt et al. (2024); Rodriguez et al. (2024). Fault-tolerant operation of logical qubits allows one to scale up the underlying QEC codes to larger distances, and thereby systematically suppress associated logical error rates. Recent experiments have demonstrated the suppression of logical error rates expected below threshold AI (2024); Google Quantum AI (2023); da Silva et al. (2024); Gupta et al. (2024); Hong et al. (2024); Sivak et al. (2023).
@@
The exact threshold value depends on the QEC code and decoding procedure used for recovering the logical information. Among all decoding strategies, Maximum Likelihood Decoding (MLD) stands out because it has the highest possible threshold, dubbed optimal threshold. Then, a straightforward approach for obtaining optimal thresholds involves simulating QEC by sampling over all possible error processes and then computing the corrective operation using MLD. The latter requires, in general, to solve an exponentially hard problem Iyer and Poulin (2015); Fuentes et al. (2021) every time an error is corrected, thus severely limiting the size of the codes one can simulate. However, recently it has been pointed out that the coherent information (CI) Schumacher and Nielsen (1996) of the mixed state associated to the noisy qubits forming the QEC code shows a discontinuity at the optimal threshold Fan et al. (2024); Colmenarez et al. (2024). In fact, the CI captures the information left in the noisy mixed-state that can, in principle, be recovered Schumacher and Nielsen (1996). Therefore, the CI provides a practical, generic tool for estimating optimal thresholds without resorting to numerical simulation of noisy QEC cycles and performing MLD.
@@
Once one or more physical qubits are lost or erased, the corrective operation is equivalent to trying to retrieve the logical qubit from the remaining physical qubits. Several QEC schemes for the correction of erasure errors have been designed Grassl et al. (1997); Amaro et al. (2020); Auger et al. (2017); Delfosse and Nickerson (2021); Goto and Uchikawa (2014); McLauchlan et al. (2024); Morley-Short et al. (2019); Nagayama et al. (2017); Siegel et al. (2023); Steinberg et al. (2024); Vodola et al. (2022), including first experimental demonstrations of deterministic correction of qubit loss Stricker et al. (2020); Reichardt et al. (2024). Let us note that optimal decoding of QEC codes in the presence of qubit loss is a fundamentally different problem compared to computational errors. For instance, in the case of the three-dimensional surface code, there is a linear-time maximum likelihood decoder (MLD) under qubit loss Delfosse and Zémor (2020), whereas efficient optimal decoding for computational errors has been demonstrated only for two-dimensional surface codes Bravyi et al. (2014); Chubb (2021). Another difference concerns the minimum number of physical qubits required to correct a single error on a qubit, which is four for qubit losses Grassl et al. (1997) and five for an arbitrary computational error Laflamme et al. (1996). Furthermore, optimal thresholds of QEC codes under erasure errors are usually higher than for computational errors Stace et al. (2009). In fact, given the positive trade-off in correcting qubit losses instead of computational errors, it has been proposed that erasure conversion, i.e., converting computational errors into qubit losses, could significantly improve the overall performance of error correction protocols Wu et al. (2022); Kang et al. (2023); Sahay et al. (2023); Quinn et al. (2024).
@@
On the other hand, we numerically compute the CI for different values of the error probability and find that the finite-size crossings of the CI in small distance codes yield very accurate approximations to the optimal error thresholds compared to rigorous solutions of the corresponding statistical mechanics models Stace et al. (2009); Stace and Barrett (2010); Ohzeki (2012). This provides further evidence of the utility of the CI in estimating optimal thresholds of QEC codes from small code instances Colmenarez et al. (2024). Let us note that in this work we do not intend to solve the statistical mechanics models and pinpoint their phase transition, as done in several other works Dennis et al. (2002); Chubb and Flammia (2021). Instead, we use the models as a mean to numerically compute the CI and extract optimal thresholds directly from it. To showcase the application of our method beyond topological codes, we perform a similar study on a low-density parity-check (LDPC) code—the lift-connected surface code Old et al. (2024). We find optimal thresholds for erasure errors only and derive statistical mechanics mappings for both computational and erasure errors. To the best of our knowledge, the models derived have not been studied previously in the literature. Our methods allow for a direct derivation of these mappings for LDPC codes and open the door to new families of statistical mechanics models beyond the realm of topological codes.
The structure of the paper is as follows. In Sec. II a summary of the main results of this work is presented. In Sec. III the main concepts used are explained in detail. In Sec. IV we present the main derivations of our work in detail. In Sec. V we show the numerical calculations of CI for surface and color codes. In Sec. VI, we apply the same methodology to the lift-connected surface code and present numerical results. Finally, concluding remarks are presented in Sec. VII.
@@ II.2 Coherent information mapping of CSS codes under depolarizing noise
Table 1: Pauli error probabilities for the Pauli channel of Eq. (3) on each site $\ell$ in term of the random variables $\eta^{x}_{\ell}$ and $\eta^{z}_{\ell}$. $C^{(1)}$ shows the connection to the error chains in Sec. IV.
@@
$H[\{\eta^{x}\},\{\eta^{z}\}]=\sum_{\ell}\eta^{x}_{\ell}\left(J_{x}-\dfrac{J_{1}}{2}\right)\dfrac{P^{X}_{\ell}}{2}+$
$\eta^{z}_{\ell}\left(J_{z}-\dfrac{J_{1}}{2}\right)\dfrac{P^{Z}_{\ell}}{2}+\eta^{z}_{\ell}\eta^{x}_{\ell}J_{1}\dfrac{P^{Z}_{\ell}P^{X}_{\ell}}{4}.$ (5)
@@ II.3 Coherent information mapping of CSS codes under depolarizing and erasure errors
The last equation shows the processes that reduce the CI: (i) qubit losses can degrade the logical qubit to a classical bit or completely destroy it, and (ii) the cost of computational errors and their interplay with qubit losses is modeled by the partition functions $\mathcal{Z}_{D_{l}}$, with missing bonds at the erased qubit positions. We can reason about these two processes as follows: erasure errors weaken the ability to protect quantum information by rendering some stabilizers ill-defined and effectively reducing the number of degrees of freedom used for error detection and correction. However, in the absence of computational errors, this does not affect the logical qubit unless at least one logical operator becomes ill-defined. In the presence of computational errors, there are fewer operators that commute with the well-defined stabilizers, but the partition functions $\mathcal{Z}_{D,l}$ still count the number of operators that commute with the well-defined stabilizers. Note that $D_{l}$ now denotes the set of operators that anti-commute with the well-defined logical operators. For details, see Sec. IV.3.
@@ II.4.2 Optimal thresholds from numerical calculation of CI
II.5 Lift-connected surface code
Since the CI methodology developed in our work applies directly to low-density parity-check (LDPC) codes, we apply this machinery to the lift-connected surface code (LCS) Old et al. (2024). This code features a growing number of logical qubits, which is clearly distinct from topological codes. As shown in Fig. 11, the LCS code can be viewed as a periodic array of interconnected surface code sheets. Each surface code stabilizer is extended over neighboring surface code sheets. Despite the stabilizers having 3D-local connectivity, the structure of the logical operators does not resemble that of a topological code. We numerically compute the coherent information (CI) for erasures only and find a $50\%$ threshold, along with a critical exponent that neither matches the 2D nor the 3D percolation transition. We then study the interplay between computational and erasure errors by deriving the corresponding statistical mechanics mapping and computing the CI numerically. The Hamiltonian of the model reads:
$H_{LCS}=\sum_{\langle i,j\rangle,\langle n,m\rangle,q}\eta^{x}_{ij}\left(J_{x}-\dfrac{J_{1}}{2}\right)\dfrac{\sigma_{i,q}\sigma_{j,q}\sigma_{i,q+1}}{2}$
$+\eta^{z}_{nm}\left(J_{z}-\dfrac{J_{1}}{2}\right)\dfrac{\tau_{n,q}\tau_{m,q}\tau_{m,q+1}}{2}$
$+\eta^{z}_{nm}\eta^{x}_{ij}J_{1}\dfrac{\sigma_{i,q}\sigma_{j,q}\sigma_{i,q+1}\tau_{n,q}\tau_{m,q}\tau_{m,q+1}}{4}.$ (12)
The indices $\langle i,j\rangle$ and $\langle n,m\rangle$ denote qubit positions on the $q$-th surface code sheet. The two-spin interactions become three-spin interactions by coupling one spin from the same “plaquette” or “star” to a spin from the next surface code sheet. One instance of this model is depicted in Fig. 12. This new family of statistical mechanics models is 3D-local and exhibits directionality in the interactions (see Appendix C for further details). Furthermore, we perform a pseudo-threshold analysis for bit/phase-flip and depolarizing noise combined with erasure errors. We find thresholds comparable to those of the standard surface code, despite the growing number of logical qubits. We conclude that the LCS code offers similar levels of protection as the surface code, with the added benefit of supporting a growing number of logical qubits.
@@ III.1 Stabilizer codes
In particular we are interested in codes for which the stabilizer generators $g_{i}$ are composed of either $X$ or $Z$ Pauli operators, i.e. Calderbank-Shor-Steane (CSS) codes Calderbank and Shor (1996); Steane (1997). This class of codes allows the logical operators $\{O^{x}_{i},O^{z}_{i}\}$ to be made of only $Z$ or $X$ Pauli operators. Within the class of CSS codes we are interested in topological codes Kitaev (2006). They are defined on $D-$dimensional lattices and encode a fixed number $k$ of logical qubits, i.e. $k$ is independent of $n$ and only dependent on the topological properties of the manifold the lattice is embedded in. The stabilizer generators $\{g_{i}\}$ are defined on unit cells of the lattice and each qubit is shared by a constant number of stabilizer generators. In this work we focus on two-dimensional toric/surface and color code Kitaev (1997); Dennis et al. (2002); Bombin and Martin-Delgado (2007a) which are amongst today’s leading contenders for the realization of scalable fault-tolerant quantum computing AI (2024); Google Quantum AI (2023, 2021); Krinner et al. (2022); Zhao et al. (2022); Andersen et al. (2020); Postler et al. (2022); Ryan-Anderson et al. (2022); da Silva et al. (2024); Berthusen et al. (2024); Ryan-Anderson et al. (2021, 2024); Pogorelov et al. (2024); Postler et al. (2024); Huang et al. (2024a); Bluvstein et al. (2024); Lacroix et al. (2024); Rodriguez et al. (2024). We refer to Appendix B for definitions and details of the topological codes studied in this work. Besides topological codes, we study the lift-connected surface code Old et al. (2024), which belongs to the class of CSS low-density parity-check (LDPC) codes. This code is part of the hypergraph product code family and has the distinctive feature that it can be viewed as an array of interconnected surface codes. When the code parameters are appropriately chosen, the rate and distance of the code exceed those of disconnected surface codes.
@@ III.2.1 Erasure channel
There is a probability $e$ that the state $\rho$ is replaced by a maximally mixed state. We are interested in the case when we know the qubit is lost due to the action of the channel. Therefore, a more convenient way to describe this process is to define a CPTP map that attaches one ancilla bit to each Kraus operator of the channel such that
$\mathcal{E}_{1}(\rho)=(1-e)\rho\otimes|\uparrow\rangle\langle\uparrow|+e\frac{\mathbb{1}}{2}\otimes|\downarrow\rangle\langle\downarrow|$ (14)
[…]
$\mathcal{E}_{n}(\rho)=\sum_{l=1}^{2^{n}}P(l)\frac{1}{2^{m}}\mathbb{1}_{A_{l}}\otimes\operatorname{Tr}_{A_{l}}(\rho)\otimes|a_{l}\rangle.$ (15)
Here $A_{l}$ denotes the set of erased qubits and $l=1,2,..,2^{n}$ runs over all possible configurations of erased qubits, $|a_{l}\rangle$ is the configuration of the ancilla bits guaranteed to satisfy $\langle a_{l^{\prime}}|a_{l}\rangle=\delta_{ll^{\prime}}$ and $m\equiv m(l)$ is the number of qubits in the erased region $A_{l}$. When erasures are independent on each qubit, e.g. all qubits experience the same erasure operation given by Eq. (14) with the same probability $e$, then $P(l)=(1-e)^{1-m}e^{m}$. In general a distance-$d$ QEC code can correct all erasure configurations with $m<d$. All erasure configurations for which at least one representative of each logical operator remains well-defined is fully correctable.